OpenStudy (staceyg):

Write an equation in point-slope form for the line through the given point with the given slope. (8,3); m=6

OpenStudy (staceyg):

y+3=6(x-8) y-3=6(x-8) y-3=6(x+8) y+3=6x+8

OpenStudy (texaschic101):

point-slope formula : y - y1 = m(x - x1) slope(m) = 6 (8,3) x1 = 8 and y1 = 3 now we sub y - 3 = 6(x - 8) any questions ?

OpenStudy (staceyg):

Oh wow that was easier than I thought it would be. Thanks, no questions. :) I just need to remember the formulas.

OpenStudy (texaschic101):

lol....there is so many formulas.....I don't remember them all

OpenStudy (staceyg):

So basically for every question like this one, you just take the formula and replace the variables with the info from the question to get the answer?

OpenStudy (texaschic101):

yep...if they want point slope form. Just be careful for the signs. For instance, if your points were (-8,3) and the slope of 6 your answer would be : y - y1 = m(x - x1) y - 3 = 6(x - (-8) y - 3 = 6(x + 8) understand ?

OpenStudy (staceyg):

Yes, I understand. Thanks for all the help :)

OpenStudy (texaschic101):

anytime :)